OpenStudy (anonymous):

The f(x)=(1/2)x^3-2x+3 Use the limit definition of area to find the area of the region betweeen the graph of the function and the x-axis on the given interval, [1,3]

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

do you have a hint or notes on this

OpenStudy (watchmath):

Divide the interval [1,3] into n intervals each of length 2/n. Then the area is given by \(\lim_{n\to\infty}\sum_{i=1}^nf\left(1+\frac{2i}{n}\right)\frac{2}{n}\)